OpenStudy (anonymous):

What is the simplified form of 5x+4/x+7-2x+1/x+7 ?

hero (hero):

\[\frac{5x+4}{x+7}-\frac{2x+1}{x+7}\]

hero (hero):

\[\frac{5x+4 - (2x + 1)}{x+7}\]

OpenStudy (anonymous):

Thankk Youuuu :D

hero (hero):

That is not the full solution